The firm is an all-equity firm with assets worth $350 million and 100 million shares outstanding. It plans to borrow $100 million and use these funds to repurchase shares. The firm’s marginal corporate tax is 21%, and it plans to keep its outstanding debt equal to $100 million permanently. If the firm manages to repurchase shares at $4 per share, what is the per share value of equity for the leveraged firm?

A) $2.71 per share

B) $3.5 per share

C) $3.61 per share

D) $3.71 per share

E) $4 per share
Answer #1

Total equity value = Value of unlevered firm+ tax(debt) - debt

Total equity value = 350,000,000 + 0.21(100,000,000) - 100,000,000

Total equity value = 271,000,000

Purchased shares = 100,000,000 / 4 = 25,000,000

Outstanding shares after repurchase = 100,000,000 - 25,000,000 = 75,000,000

Share price = 271,000,000 / 75,000,000

Share price = $3.61 per share